The function f, defined by f(x)= |x-2| is differentiable in [0,1].

True or false with full explanation




1
Expert's answer
2021-05-31T00:10:27-0400

Let us show that the function "f", defined by "f(x)= |x-2|" is differentiable in "[0,1]". If "x-2<0", that is "x<2", then "f(x)=-(x-2)=-x+2." In particular, for "x\in[0,1]" we have that "f(x)=-x+2." Since the elementary linear function is differentiable in each point, we conclude that the function "f" is differentiable in "[0,1]".